# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v11

@@ 9 Regressive Generation of Sin and Cos
Figure 57: The uncertainties (y-axis) of $\sin(\pi j/2^{18})$ as a function of $j$ (x-axis) and $L$ using either Quart or regressive sine function (legend).
Figure 58: Uncertainties, value errors (left y-axis), and error deviation (right y-axis) for $\sin(x)^{2}+\cos(x)^{2}-1,x\in[0,\pi/4]$ as a function of regression order (x-axis) when compared with Quart sine function (legend)
@@
Formula (9.2) is not suitable to compute $\sin(x)$ when $x\rightarrow 0$ because it fails similarly as catastrophic cancellation, obtaining excessive uncertainty. In Figure 57, Quart sine has uncertainties proportional to $\sin(\pi j/2^{L})$, while regression sine has its uncertainties increase when $x\rightarrow 0$.
Figure 58 tests Quart and Regression sin/cos for $\sin(x)^{2}+\cos(x)^{2}-1=0$. In Figure 58, the result uncertainties are nearly a constant independent of the regression order $L$ when $L\geq 4$ for both Regression sine and Quart sine. It also shows that the value errors for both are quite comparable, resulting in the error deviation of the Regression sin/cos closer to $1$.
@@ 10.1 Summary of Statistical Taylor Expansion
The starting point of statistical Taylor expansion is the assumption that all input variables have uncorrelated uncertainties. This is considered a very reasonable statistical requirement on the input data [1]. Once this assumption is met, statistical Taylor expansion quantifies uncertainty as the deviation of the value errors. It can track the variable dependence in the intermediate steps using standard statistical methods. While it eliminates the dependency problem, it requires a more rigorous process to determine resulting mean and deviation for an analytic expression. In addition, it can reject calculations on unified statistical and mathematical grounds. It also incorporates sample count into statistical analyses.
@@ 10.2 Summary of Variance Arithmetic
Variance arithmetic simplifies statistical Taylor expansion by assuming that all inputs are Gaussian distributed and with enough sample counts for ideal variance. Ill-formed problems can invalidate resulting variance in several ways: non-convergent, unstable, negative, infinite, and unreliable.
The presence of ideal coverage is a necessary condition for a numerical algorithm in variance arithmetic to be considered correct. Ideal coverage also defines the algorithm’s optimal applicable range for the algorithm. Under ideal coverage, the calculated uncertainty equals the deviation of the value errors, and the normalized errors follow either a Normal or Delta distribution depending on the context.
Variance arithmetic can provide proper coverage for floating-point rounding errors.
The applicability of variance arithmetic has been demonstrated in diverse scenarios, including analytic calculations, progressive computation, regressive generalization, polynomial expansion, statistical sampling, and transformations.
@@ 10.3 Improvements Needed
This paper presents variance arithmetic which is still in its early developmental stage. From a theoretical perspective, several important questions remain to be addressed.
Library mathematical functions should be recalculated using variance arithmetic, so that each output value is accompanied by its corresponding uncertainty. Without this, the value errors in the library functions can produce unpredictable and potentially large effects on numerical results.
@@
Determining the uncertainty upper bounds for each analytic expression analytically is important. Under current numerical framework, the measured upper bounds in Figure 7 for $(x+\delta x)^{c}$, and Figure 9 for $\sin(x+\delta x)$ may change with implementation such as between 32-bit and 64-bit floating-point computations.
@@
The performance of variance arithmetic has to be improved. The fundamental formulas for statistical Taylor expansion, Formula (2.6), (2.3), (2.9), and (2.3) contains large number of independent summations, making them excellent candidates for parallel processing, such as direct hardware implementation. The procedural nature of these formulas allows statistical Taylor expansion to be implemented directly in hardware.
When an analytic expression undergoes statistical Taylor expansion, the result expression may be very complex, such as in matrix inversion. However, modern symbolic computation tools such as SymPy can greatly facilitate these calculations. This suggests that it may be time to shift from purely numerical programming toward analytic programming for problems with inherently analytic formulations.
As an enhancement to dependency tracing, source tracing identifies each input’s contribution to the overall result uncertainty. This capability can help engineers pinpoint the primary drivers of measurement inaccuracy and thus guide targeted improvements in data acquisition and processing strategies.
A key open question is whether variance arithmetic can be adapted to provide ideal coverage for floating-point rounding errors. Once the mechanism for rounding error is thoroughly understood, variance arithmetic with ideal coverage to floating-point rounding errors is quite valuable because many theoretical calculations lack explicit input uncertainty.
Because traditional numerical approaches are based on floating-point arithmetic, they need to be reexamined or even reinvented in variance arithmetic. For example, most conventional numerical algorithms aim to select optimal computational paths, whereas variance arithmetic fundamentally rejects conceptually all path-dependent calculations. Reconciling these two paradigms may present a significant challenge.
Establishing theoretical foundation to apply statistical Taylor expansion in the absence of analytic solution, or with only limited low-order numerical derivatives, such as in solving differential equations, remains an important research direction.
@@ 10.4 Acknowledgments
As an independent researcher without institutional affiliation, the author expresses deep gratitude for the encouragement and valuable discussions provided by Dr. Zhong Zhong (Brookhaven National Laboratory), Prof Weigang Qiu (Hunter College). Special thanks are extended to the organizers of AMCS 2005, particularly Prof. Hamid R. Arabnia (University of Georgia), and organizers of NKS Mathematica Forum 2007. Prof Dongfeng Wu (Louisville University) is acknowledged for insightful guidance on statistical topics. The author is also indebted to the editors and reviewers of Reliable Computing for their substantial assistance in shaping and accepting a previous version of this work, with special recognition to Managing Editor Prof. Rolph Baker Kearfott.

### v12

@@ 10 Regressive Generation of Sin and Cos
Figure 60: The uncertainties (y-axis) of $\sin(\pi j/2^{18})$ versus $j$ (x-axis) and $L$ using either Quart or regressive sine function (legend).
Figure 61: Uncertainties, value errors (left y-axis), and error deviation (right y-axis) for $\sin(x)^{2}+\cos(x)^{2}-1,x\in[0,\pi/4]$ versus regression order (x-axis) when compared with the Quart sine function (legend)
@@
Formula (10.2) is not suitable for computing $\sin(x)$ as $x\rightarrow 0$ because it suffers from behavior analogous to catastrophic cancellation, resulting in excessive uncertainty. In Figure 60, the Quart sine function exhibits uncertainties proportional to $\sin(\frac{\pi}{2^{L}})$, whereas the regression sine function shows increasing uncertainties as $x\rightarrow 0$. Unlike catastrophic cancellation in floating-point arithmetic, variance arithmetic uses very coarse precision to signal that the regression algorithm is unfit to compute $\sin(x\rightarrow 0)$.
Figure 61 evaluates the Quart and Regression sine functions using $\sin(x)^{2}+\cos(x)^{2}-1=0$. The figure shows that the resulting uncertainties remain nearly constant and independent of the regression order $L$ when $L\geq 4$ for both the Regression sine and the Quart sine functions. It also demonstrates that the value errors for both are comparable. The Regression sine function has error deviations closer to $1$.
@@ 11.1 Summary of Statistical Taylor Expansion
The starting point of statistical Taylor expansion is the assumption that all input variables obeys uncorrelated uncertainties. This is generally regarded as a reasonable and practical statistical requirement for input data [1]. Once this assumption is satisfied, statistical Taylor expansion quantifies uncertainty as the deviation of the value errors and traces track the variable dependencies in intermediate steps using standard statistical methods. While this approach eliminates the dependency problem, it also requires a more rigorous process to determine the resulting mean and deviation for any analytic expression. Furthermore, it can reject invalid calculations on unified statistical and mathematical grounds and explicitly incorporates sample counts and the corresponding underlying uncertainty distributions into the statistical analysis.
@@ 11.2 Summary of Variance Arithmetic
Variance arithmetic simplifies statistical Taylor expansion by assuming that the sample counts for all inputs are sufficiently large to achieve ideal variance. Ill-formed problems can invalidate the resulting variance in several ways by producing non-convergent, unstable, negative, infinite, or unreliable results.
The presence of ideal coverage is a necessary condition for a numerical algorithm in variance arithmetic to be considered correct. Ideal coverage also defines the algorithm’s optimal applicable range. Under ideal coverage, the calculated uncertainty equals the deviation of the value errors, and the normalized errors follow either a Normal or Delta distribution depending on the context.
Variance arithmetic provides proper coverage for floating-point rounding errors.
The applicability of variance arithmetic has been demonstrated across a wide range of computational scenarios, including analytic calculations, progressive computation, regressive generalization, polynomial expansion, statistical sampling, and transformations.
@@ 11.3 Improvements Needed
This paper presents variance arithmetic which is still in its early stage of development. From a theoretical standpoint, several important questions still need to be addressed.
Library mathematical functions should be recalculated using variance arithmetic, so that each output value is accompanied by its corresponding uncertainty. Without this refinement, the value errors in the library functions can produce unpredictable and potentially significant effects on numerical results. How to compute and validate these new uncertainty-bearing library functions will need innovation. For example, the uncertainty of a 64-bit value can not directly determined by its difference with the corresponding 128-bit value,
@@
Determining the uncertainty upper bounds for each analytic expression through analytic methods is an important next step. These results can validate variance arithmetic, and provide performance enhancement to variance arithmetic when the convergence of an analytic expression has to be deduced numerically from the full expansion each time.
The performance of variance arithmetic must be improved for broader practical adoption. The fundamental formulas of statistical Taylor expansion, Formula (2.9), (2.3), (2.12), and (2.3) contain a large number of independent summations, making them excellent candidates for parallel processing. Moreover, the inherently procedural nature of these formulas allows statistical Taylor expansion to be implemented efficiently at the hardware level.
When an analytic expression undergoes statistical Taylor expansion, the resulting expression can become highly complex, as in the case of matrix inversion. However, modern symbolic computation tools such as SymPy can greatly facilitate these calculations. This observation suggests that it may be time to shift from purely numerical programming toward analytic programming, particularly for problems that possess inherently analytic formulations.
As an enhancement to dependency tracing, source tracing identifies each input’s contribution to the overall result uncertainty. This capability enables engineers to pinpoint the primary sources of measurement inaccuracy and in turn guide targeted improvements in data acquisition and processing strategies.
A key open question is whether variance arithmetic can be adapted to achieve ideal coverage for floating-point rounding errors. Once the mechanism of rounding error is thoroughly understood, developing variance arithmetic with ideal coverage for such errors would be quite valuable because many theoretical calculations lack explicit input uncertainties.
Because traditional numerical approaches are based on floating-point arithmetic, they must be reexamined or even reinvented within the framework of variance arithmetic. For instance, most conventional numerical algorithms aim to identify optimal computational paths, whereas variance arithmetic conceptually rejects all path-dependent calculations. Reconciling these two paradigms may present a significant and ongoing challenge.
Establishing theoretical foundation for applying statistical Taylor expansion in the absence of a closed-form analytic solution, or when only limited low-order numerical derivatives are available, as in solving differential equations, remains an important direction for future research.
@@ 11.4 Acknowledgments
As an independent researcher without institutional affiliation, the author expresses sincere gratitude to Dr. Zhong Zhong (Brookhaven National Laboratory) and Prof Weigang Qiu (Hunter College) for their encouragement and valuable discussions. Special thanks are extended to the organizers of AMCS 2005, particularly Prof. Hamid R. Arabnia (University of Georgia), and to the organizers of the NKS Mathematica Forum 2007. The author also gratefully acknowledges Prof Dongfeng Wu (Louisville University) for her insightful guidance on statistical topics. Finally, heartfelt appreciation is extended to the editors and reviewers of Reliable Computing for their substantial assistance in shaping and accepting an earlier version of this work, with special recognition to Managing Editor Prof. Rolph Baker Kearfott.